OpenStudy (anonymous):

Write the equation of the circle with center (-6, -4) and containing the point (-2, -1). Use the ^ key for the exponents. Write your answer as the example: (x-2)^2+(y+4)^2=64

OpenStudy (anonymous):

(x+6)^2 +(y+4)^2 = 89

OpenStudy (dumbcow):

solve for r^2 (x+6)^2 + (y+4)^2 = r^2 x=-2, y=-1 4^2 + 3^2 = r^2 25 = r^2 (x+6)^2 + (y+4)^2 = 25

OpenStudy (anonymous):

come now samephone18, you should be able to these problems on your own now. You have been given many examples of how to do this exact same problem

OpenStudy (amistre64):

lol ... answering questions is prolly best when done in the spirit of putting out a useable answer, and not really as the objective of teaching any one individual

OpenStudy (anonymous):

oh well, I disagree. I think if we can show them how to do one problem clearly, then they will have the tools necessary to do other problems and variations. It is more important to learn to think about how to approach problems rather than actually doing the problem. That is my opinion of course :)